OpenStudy (anonymous):

find the area of the rectangle if the length 2x^3 and the width is 7x^3

OpenStudy (anonymous):

what is the formula for rectangle area ?

OpenStudy (anonymous):

A=lw

OpenStudy (anonymous):

good so we need to do Area = (7x^3) * (2x^3)

OpenStudy (anonymous):

(ax^n) * (bx^m) = abx^(n+m) can you do it now ?

OpenStudy (anonymous):

The common equation for the area of a rectangle is Area = Length x Width, right? So, you have length and width. Area = (2x^3) * (7x^3). From here it's just algebraic multiplication. Simply multiply their coefficients (the 2 and the 7) and ADD their powers (3+3). Hope that helps.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is it 14x^6

OpenStudy (anonymous):

using : (ax^n) * (bx^m) = abx^(n+m) we get: A = (7x^3) * (2x^3) = (2*7*x^(3+3))

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

awesome thanks :o)